The catapult of the aircraft carrier USS Lincoln accelerates an F/A-18 Hornet jet fighter from rest to take off speed of 173mi/h in a distance of 307ft. Assume a constant acceleration. a) Calculate the acceleration of the jet fighter in m/s2 b) Calculate the time required for the fighter to be accelerated to


1
Expert's answer
2021-03-29T16:47:38-0400

(a)

"a=\dfrac{v^2-v_0^2}{2d}=\dfrac{(77.34\ \dfrac{m}{s})^2-0}{2\cdot93.57\ m}=32\ \dfrac{m}{s^2}."

(b)

"t=\dfrac{v-v_0}{a}=\dfrac{77.34\ \dfrac{m}{s}-0}{32\ \dfrac{m}{s^2}}=2.42\ s."